OpenStudy (kanika):

Twice a no. Is 12 times more than half the no. What is the no.?

OpenStudy (mrnood):

if 'the number' is x then 'twice the number' is written 2x 'half the number' is written x/2 Can you write an equation now that puts your question into 'maths' from the words?

OpenStudy (radar):

@kanika Did you make up this problem? Or is there some sadistic jerk pulling your chain?

OpenStudy (mrnood):

@radar @kanika Whilst this question does not fall into the very simplest of arithmetic if you follow my instrucyion you will have an equation. At first glance the answer may seem difficult - but there IS one value of x for which the equation works. Write the equation first - then post if you need help to solve it...

OpenStudy (kanika):

Sure

OpenStudy (radar):

Let x = the "number then x/2 = 1/2 the number 2x = twice the number. 2x = 12 + x/2 " twice a number is 12 times more than half a number" or is it 2x = (12) x/2 " twice a number is 12 times more than half a number" or is it neither ?? define "12 times more than" Well the last equation goes nowhere x=(6x)/2 or x = 3x Nonsensical answer. The first one results in x=8. The problem needs some clarification!

OpenStudy (radar):

@MrNood please enlighten us on the syntax needed here to provide the clarification.

OpenStudy (mrnood):

2x = 12 x/2 2x = 6x there is one solution

OpenStudy (radar):

Well I guess that's it............x = 0

OpenStudy (kanika):

I hv got d answer dat is 8.

OpenStudy (radar):

Or \[x = \infty \]

OpenStudy (radar):

Yes 8 if we use 2x = 12 + x/2 It is as we would say a "weasel worded question" lol

OpenStudy (mrnood):

"12 times more" is certainly not 12+x/2 Are you sure that the question is written correctly? If it were: "Twice a no. Is 12 more than half the no" We would have 2x = 12 +x/2 ->x=8